OpenStudy (luhvon):

What is 2 3/4 as a decimal?

OpenStudy (mariamikayla):

2 3/4= 11/4= 2,75

OpenStudy (luhvon):

thank u

OpenStudy (mariamikayla):

You're welcome :D